Question:


The organic compound below is abbreviated Tris and is basic. The conjugate acid (TrisITT has a pKa of 8.10.


TrisH+ + H20 = Tris + H30 +


Tris is commonly used as a component in buffer solutions to study biochemical reactions.


a) A buffer solution is prepared by mixing 12.5 mL of 0.100 mol L- 1 HCl with 25.0 mL of0.100 mol L- 1 Tris. What is the pH of this buffer? Write all the equations used and explain your steps.


b) The buffer was used to study an enzyme-catalyzed reaction in which 0.00020 mol of H30+ was produced. What is the pH of the solution after the acid reacted with the buffer? Write all the equations used and explain your steps.


1
Expert's answer
2022-03-27T10:30:56-0400

(a) pKa for conjugate acid = 8.10

volume of HCl = 12.5 mL, Molarity= 0.1M

volume of Tris = 25 mL, Molarity = 0.1 M

pH= pKa+ "\frac{log (Tris)}{HCl}"


"pH= 8.10 +\frac{log (0.1)}{0.1}= 8.10"


(b) mole of H3O+ produced = 0.0002 mol


"pH= 8.10 +\frac{log (0.1)}{0.0002}= 7.90"